OpenStudy (anonymous):

What is the solution of |x-4| <= -5 ?

OpenStudy (anonymous):

Do you mean that to be a greater than or equal to sign?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

x - 4 = -5 x = -1 x - 4 = 5 x=9. x = -1,9 x|-1<x<9 Technically, this problem makes no sense since it's an absolute value sign. Which means the outcome can't equal a negative number.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The following is the graph of -2x + y = 7. True or False on the graph.